### v2

@@ Abstract
Visual recognition relies on understanding both the semantics of image tokens and the complex interactions among them. Mainstream self-attention methods, while effective at modeling global pair-wise relations, fail to capture high-order associations inherent in real-world scenes and often suffer from redundant computation. Hypergraphs extend conventional graphs by modeling high-order interactions and offer a promising framework for addressing these limitations. However, existing hypergraph neural networks typically rely on static and hard hyperedge assignments, leading to excessive and redundant hyperedges with hard binary vertex memberships that overlook the continuity of visual semantics. To overcome these issues, we present Soft Hypergraph Neural Networks (SoftHGNNs), which extend the methodology of hypergraph computation, to make it truly efficient and versatile in visual recognition tasks. Our framework introduces the concept of soft hyperedges, where each vertex is associated with hyperedges via continuous participation weights rather than hard binary assignments. This dynamic and differentiable association is achieved by using the learnable hyperedge prototype. Through similarity measurements between token features and the prototype, the model generates semantically rich soft hyperedges. By using these soft hyperedges as the medium for message aggregation and dissemination, SoftHGNN models various high-order semantic relationships within visual features. To further enhance efficiency when scaling up the number of soft hyperedges, we incorporate a sparse hyperedge selection mechanism that activates only the top-$k$ important hyperedges, along with a load-balancing regularizer to ensure adequate and balanced hyperedge utilization. Experimental results across three tasks on five datasets demonstrate that SoftHGNN efficiently captures high-order associations in visual scenes, achieving significant performance improvements. The source code of our framework is available at: https://github.com/Mengqi-Lei/SoftHGNN.
@@ 1 Introduction
In recent years, significant advancements in visual recognition have primarily been driven by deep neural network-based models, especially Convolutional Neural Networks (CNNs) and Transformer models [1, 2, 3]. CNNs excel at capturing local spatial patterns, whereas Transformer models, built upon self-attention mechanisms, effectively model global relationships among image tokens. Current mainstream paradigms, such as Vision Transformer (ViT) [4] and its variants [5, 6, 7, 8, 9], typically process images into sequences of visual tokens, emphasizing the modeling of pair-wise relations between these tokens.
@@
Despite achieving remarkable success, vision models based on self-attention mechanisms [10, 4] still face significant challenges. Firstly, self-attention mechanisms inherently construct a fully connected semantic graph that primarily focuses on pair-wise relationships among tokens, making it difficult to effectively represent complex high-order semantic associations commonly present in real-world visual scenes. For instance, the simple scene of “a person holding a tennis racket and hitting the ball” involves the interaction among multiple entities, such as the connection among the person, the tennis racket, and the sports ball. Such high-order associations cannot be adequately captured by merely modeling pair-wise relations. Secondly, Transformers typically employ dense global self-attention, leading to substantial redundant computations, high computational costs, and difficulties in model convergence.
@@
To alleviate these issues, hypergraph, as an extension of traditional graph structure, has received considerable attention due to its intrinsic capability to explicitly model high-order associations [11, 12]. Hypergraph Neural Networks (HGNNs) explicitly represent relationships among multiple vertices via hyperedges, making them promising for capturing complex semantic interactions in visual data [13, 14]. Recent studies, such as Hyper-YOLO [15], have initially validated the effectiveness of HGNNs in various vision tasks. However, existing HGNN methods, originally designed for network-type data, encounter critical issues when applied to visual tasks: 1) Redundant hyperedges. Traditional hypergraph constructions typically employ methods like $k$-Nearest Neighbors or $\epsilon$-ball criteria, which generate one hyperedge centered around each vertex. In visual tasks, given the large number of vertices (typically $B\times H\times W$, where $B$ is batch size), this approach often results in an excessively large number of hyperedges, far exceeding the actual number of meaningful high-order associations. Consequently, this leads to significant computational redundancy and drastically reduces model efficiency. 2) Limitations of hard hyperedges. Existing HGNNs typically rely on hard vertex-to-hyperedge associations, meaning a vertex is either completely included in a hyperedge or entirely excluded from it. This hard binary assignment overlooks the inherent continuity and ambiguity of visual semantics. For example, in partially occluded scenarios, regions of occluded objects should ideally participate in relevant hyperedges with lower weights, rather than being completely excluded. Moreover, hard hyperedge designs often cause vertex redundancy or incomplete coverage, introducing substantial noise and severely impairing model performance.
@@
To overcome these bottlenecks, we propose the Soft Hypergraph Neural Network (SoftHGNN), which extends the methodology of hypergraph computation, making it genuinely general and efficient for visual recognition tasks. The core innovation is transitioning from static and hard hyperedge assignments to a dynamic and feature-driven approach with a differentiable soft vertex-participation mechanism. Specifically, different from traditional hyperedge construction methods based on geometric distances or feature similarities, SoftHGNN employs a set of learnable hyperedge prototype vectors. By measuring the similarity between visual tokens and these prototype vectors, the model dynamically generates semantically rich “soft hyperedges”. Each hyperedge softly connects all vertices with learnable participation weights, thus adaptively modeling abstract high-order semantic relationships within visual features. Furthermore, SoftHGNN only needs to maintain a constant-scale set of soft hyperedges, significantly enhancing computational efficiency and effectiveness compared to traditional HGNNs. Moreover, the proposed SoftHGNN is a plug-and-play module that can be easily integrated into the model of any visual recognition task to fill the gap in the ability of existing methods to model high-order visual semantic associations, thereby bringing about significant performance improvements.
@@
To validate the generalization and effectiveness of our proposed method, extensive experiments were conducted on five mainstream datasets, namely CIFAR-10 [16], CIFAR-100 [16], ShanghaiTech Part-A [17], ShanghaiTech Part-B [17], and MS COCO [18]. These experiments cover three representative visual recognition tasks: image classification, crowd counting, and object detection. Comprehensive results demonstrate that the proposed SoftHGNN series can accurately and efficiently capture high-order semantic relationships in visual scenes, achieving significant performance improvements.
@@
  • We propose the Soft Hypergraph Neural Network (SoftHGNN) for general visual recognition. By introducing a feature-driven soft hyperedge mechanism, our model adaptively captures abstract high-order semantic relationships within visual features.
@@
  • We conducted extensive experiments on five datasets across three mainstream visual recognition tasks. Experimental results demonstrate that our method accurately and efficiently captures high-order semantic associations in visual scenes, resulting in significant performance improvements.
@@ 2 Related Work
The hypergraph is a kind of structure that naturally represents groups and high-order relationships within a system. Formally, it is an extension of a conventional graph, in which a single hyperedge can connect an arbitrary number of vertices. This structure is particularly well-suited for describing group interactions [11]. For example, in a co-authorship network, a paper can be considered a hyperedge that connects all the authors involved. Similar high-order relationships are widely present in fields including biology and neuroscience [12, 19].
@@
Hypergraph Neural Networks (HGNNs) incorporate neural networks into hypergraph representation learning. HGNNs leverage hypergraph convolution, a two‑step message‑passing scheme that first aggregates vertex features into hyperedges and then redistributes them back to vertices, to capture high‑order correlations [11, 14, 20]. Compared to traditional Graph Neural Networks (GNNs), this high-order information aggregation mechanism significantly enhances the representational power of HGNNs[11, 21]. Feng et al. [13] proposed a spectral-domain hypergraph convolution method that demonstrated outstanding performance on high-order relational data. Building on this, Gao et al. [22] introduced a more flexible approach for modeling higher‑order relationships along with a generalized spatial hypergraph convolution framework. It extends both the adaptability of hypergraph modeling and the universality of hypergraph convolution operations. Hypergraph learning has already shown significant advantages in modeling high-order relationships across various domains such as social networks [23, 24], knowledge graphs [25], recommendation systems [26, 27], action recognition [28, 29], and bioinformatics [30, 31]. It is important to note, however, that these tasks are inherently based on non-Euclidean data with graph or hypergraph structures. Hypergraphs have not yet been widely applied to data in explicit Euclidean spaces, such as visual data like images.
@@
In recent years, research has gradually begun to explore the incorporation of hypergraphs and hypergraph neural networks into visual recognition tasks [28, 29, 32]. Feng et al. proposed Hyper-YOLO [15], which integrates a hypergraph computation module into the Neck component of the YOLO model [33] to achieve cross-level and cross-spatial hypergraph semantic aggregation. Han et al. introduced Vision HGNN [34], exploring the replacement of Transformer modules with hypergraph convolution modules. Wang et al. proposed HGFormer [35], which establishes many-to-many topological connections at the patch level of images and utilizes a hypergraph attention mechanism to learn global visual relationships. Chen et al. proposed HI2R [36], which simultaneously extracts the structural information within a category and the relationship information between categories through hypergraphs for more accurate fine-grained visual recognition.
Although research on HGNNs in visual recognition has made some progress, current approaches merely apply classical hypergraphs and HGNNs to visual recognition tasks without addressing the inherent issues of hyperedge redundancy and the hard binary assignment of hyperedges, while incurring significantly high computational costs. We will focus on discussing these issues in the next section.
@@ 3.1 Hypergraph
A hypergraph is an extension of a graph that explicitly models high-order relationships within data. Formally, a hypergraph is represented as: $\mathcal{G}=(\mathcal{V},\mathcal{E}),$ where $\mathcal{V}$ denotes the set of vertices, and $\mathcal{E}$ denotes the set of hyperedges. Unlike traditional graphs, each hyperedge in a hypergraph can simultaneously connect an arbitrary number of vertices. As a result, hypergraphs naturally capture complex high-order associations among multiple vertices. The structure of a hypergraph can be described using an incidence matrix ${H}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{E}|}$, where $H_{v,e}=1$ if $v\in e$ and else $H_{v,e}=0$.
@@ 3.2 Hypergraph Neural Networks
Despite the theoretical promise of HGNNs in capturing high-order semantic associations, several inherent limitations arise when applying them to visual recognition tasks. First, traditional hypergraph constructions typically rely on methods such as $k$-Nearest Neighbors or $\epsilon$-ball criteria, which generate one hyperedge per vertex. In visual applications, where the number of vertices is extremely large, this approach leads to a redundant explosion of hyperedges. The redundancy not only introduces substantial computational overhead but also results in many hyperedges that capture overlapping or irrelevant relationships, thereby diluting the model’s focus on truly meaningful associations. Second, the incidence matrix $H$ in the traditional hypergraph enforces a hard binary assignment: a vertex is either fully included in a hyperedge or not at all. Such a rigid formulation fails to capture the inherent continuity and ambiguity present in visual data, especially in scenarios with partial occlusions or overlapping semantic regions, thus limiting the expressiveness and robustness of the model.
@@ 3.3.1 Complexity of Hypergraph Construction
Figure 1: Overview of the Soft Hypergraph Neural Network (SoftHGNN) pipeline. (1) Soft Hyperedge Generation. A set of learnable prototype vectors is dynamically offset by a context‑aware network to form sample‑specific soft hyperedges. Then each vertex compute participation weights to build the continuous participation matrix $A$. (2) Message Passing on Soft Hypergraph. High‑order interactions are captured via a two‑stage aggregation from vertices to hyperedges ($\mathcal{V}\!\to\!\mathcal{E}$) and from hyperedges to vertices ($\mathcal{E}\!\to\!\mathcal{V}$). (3) Sparse Hyperedge Selection (optional). The sparse hyperedge selection strategy is introduced to expand the number of hyperedges while avoiding a significant increase in computational complexity.

### v3

@@ Abstract
Visual recognition relies on understanding both the semantics of image tokens and the complex interactions among them. Mainstream self-attention methods, while effective at modeling global pair-wise relations, fail to capture high-order associations inherent in real-world scenes and often suffer from redundant computation. Hypergraphs extend conventional graphs by modeling high-order interactions and offer a promising framework for addressing these limitations. However, existing hypergraph neural networks typically rely on static and hard hyperedge assignments, which lead to redundant hyperedges and overlooking the continuity of visual semantics. In this work, we present Soft Hypergraph Neural Networks (SoftHGNN), a lightweight plug-and-play hypergraph computation method for late-stage semantic reasoning in existing vision pipelines. Our SoftHGNN introduces the concept of soft hyperedges, where each vertex is associated with hyperedges via continuous and differentiable participation weights rather than hard binary assignments. These weights are produced by measuring similarities between vertex features and a small set of learnable hyperedge prototypes, yielding input-adaptive and semantically rich soft hyperedges. Using soft hyperedges as the medium for message aggregation and dissemination, SoftHGNN enriches feature representations with high-order contextual associations. To further enhance efficiency when scaling up the number of soft hyperedges, we incorporate a sparse hyperedge selection mechanism that activates only the top-$k$ important hyperedges, along with a load-balancing regularizer to ensure adequate and balanced hyperedge utilization. Experimental results across three tasks on five datasets demonstrate that SoftHGNN efficiently captures high-order associations in visual scenes, achieving significant performance improvements. The source code of our framework is available at: https://github.com/Mengqi-Lei/SoftHGNN.
@@ 1 Introduction
In recent years, significant advancements in visual recognition have primarily been driven by deep neural network-based models, especially Convolutional Neural Networks (CNNs) and Transformer models [ijcv_review, unsuper_vision_survey, vit_survey]. CNNs excel at capturing local spatial patterns, whereas Transformer models, built upon self-attention mechanisms, effectively model global relationships among image tokens. Current mainstream paradigms, such as Vision Transformer (ViT) [vit] and its variants [pvt, swin, LRT_tip, eatformer_ijcv, gridformer_ijcv], typically process images into sequences of visual tokens, emphasizing the modeling of pair-wise relations between these tokens.
@@
Despite achieving remarkable success, vision models based on self-attention mechanisms [attn, vit] still face significant challenges. Firstly, self-attention mechanisms inherently construct a fully connected semantic graph that primarily focuses on pair-wise relationships among tokens, making it difficult to effectively represent complex high-order semantic associations commonly present in real-world visual scenes. For instance, the simple scene of “a person holding a tennis racket and hitting the ball” involves the interaction among multiple entities, such as the connection among the person, the tennis racket, and the sports ball. Such high-order associations cannot be adequately captured by merely modeling pair-wise relations. Secondly, Transformers typically employ dense global self-attention, leading to substantial redundant computations, high computational costs, and it has been demonstrated [attn_collapse1, attn_collapse2, attn_collapse3] that attention collapse and dispersion may occur in deep neural networks.
@@
To alleviate these issues, hypergraph, as an extension of traditional graph structure, has received considerable attention due to its intrinsic capability to explicitly model high-order associations [hg_book, hg_survey]. Hypergraph Neural Networks (HGNNs) explicitly represent relationships among multiple vertices via hyperedges, making them promising for capturing complex semantic interactions in visual data [hgnn, hg_survey4]. Recent studies, such as Hyper-YOLO [hyperyolo], have initially validated the effectiveness of HGNNs in various vision tasks. However, existing HGNN methods, originally designed for network-type data, encounter critical issues when applied to visual tasks: 1) Redundant hyperedges. Traditional hypergraph constructions typically employ methods like $k$-Nearest Neighbors or $\epsilon$-ball criteria, which generate one hyperedge centered around each vertex. In visual tasks, given the large number of vertices (typically $B\times H\times W$, where $B$ is the batch size), this approach often generates an excessively large number of hyperedges, resulting in a highly-overlapping hyperedge set, where many hyperedges are near-duplicates and contribute little additional information while substantially increasing computation for the target task. 2) Limitations of hard hyperedges. Existing HGNNs typically rely on hard vertex-to-hyperedge associations, meaning a vertex is either completely included in a hyperedge or entirely excluded from it. This hard binary assignment overlooks the inherent continuity and ambiguity of visual semantics. For example, in partially occluded scenarios, regions of occluded objects should ideally participate in relevant hyperedges with lower weights, rather than being completely excluded. Moreover, hard hyperedge designs often cause vertex redundancy or incomplete coverage, thereby introducing noisy structures and negatively impacting model performance [hg_quality, hg_quality2, hg_quality3].
@@
To overcome these bottlenecks, we propose Soft Hypergraph Neural Networks (SoftHGNN), a lightweight plug-and-play hypergraph computation method for late-stage semantic reasoning in existing vision pipelines. The core innovation is transitioning from static and hard hyperedge assignments to a dynamic and feature-driven approach with a differentiable soft vertex-participation mechanism. Specifically, different from traditional hyperedge construction methods based on geometric distances or feature similarities, SoftHGNN employs a set of learnable hyperedge prototype vectors. By measuring the similarity between visual tokens and these prototype vectors, the model dynamically generates semantically rich “soft hyperedges”. Each hyperedge softly connects all vertices with learnable participation weights, thus adaptively modeling abstract high-order semantic relationships within visual features. Furthermore, SoftHGNN only needs to maintain a constant-scale set of soft hyperedges, significantly enhancing computational efficiency and effectiveness compared to traditional HGNNs.
@@
Our SoftHGNN can be easily integrated into models as a late-stage semantic reasoning block for a wide range of visual recognition tasks to fill the gap in their ability to model high-order visual semantic associations, thereby bringing about significant performance improvements. To validate the generalization and effectiveness of our proposed method, extensive experiments are conducted on five mainstream datasets, namely ImageNet-100 [imgnet], ImageNet-1k [imgnet], ShanghaiTech Part-A [shanghaitech], ShanghaiTech Part-B [shanghaitech], and MS COCO [mscoco]. These experiments cover three representative visual recognition tasks: image classification, crowd counting, and object detection. Comprehensive results demonstrate that the proposed SoftHGNN can accurately and efficiently capture high-order semantic relationships in visual scenes, achieving significant performance improvements.
@@
  • We propose the SoftHGNN, a plug-and-play hypergraph computation method to enhance vision models for general visual recognition. By introducing a feature-driven soft hyperedge mechanism, our model adaptively captures abstract high-order semantic relationships within visual features.
@@
  • We conduct extensive experiments on five datasets across three mainstream visual recognition tasks. Experimental results demonstrate that our method accurately and efficiently captures high-order semantic associations in visual scenes, resulting in significant performance improvements.
@@ 2.1 Hypergraph Computation
The hypergraph is a kind of structure that naturally represents groups and high-order relationships within a system. Formally, it is an extension of a conventional graph, in which a single hyperedge can connect an arbitrary number of vertices. This structure is particularly well-suited for describing group interactions [hg_book]. For example, in a co-authorship network, a paper can be considered a hyperedge that connects all the authors involved. Similar high-order relationships are widely present in fields including biology and neuroscience [hg_survey, hg_survey2].
@@
Hypergraph Neural Networks (HGNNs) incorporate neural networks into hypergraph representation learning. HGNNs leverage hypergraph convolution, a two-step message-passing scheme that first aggregates vertex features into hyperedges and then redistributes them back to vertices, to capture high-order correlations [hg_book, hg_survey4, hyper3dg]. Compared to traditional Graph Neural Networks (GNNs), this high-order information aggregation mechanism significantly enhances the representational power of HGNNs [hg_book, hg_survey3]. Feng et al. [hgnn] proposed a spectral-domain hypergraph convolution method that demonstrated outstanding performance on high-order relational data. Building on this, Gao et al. [hgnnp] introduced a more flexible approach for modeling higher-order relationships along with a generalized spatial hypergraph convolution framework, extending both the adaptability of hypergraph modeling and the universality of hypergraph convolution operations.
Recently, several works have introduced attention or Transformer-style mechanisms to parameterize vertex-hyperedge interactions. Jiang et al. [dhgnn] proposed the Dynamic Hypergraph Neural Network, which dynamically updates the hypergraph structure across layers, aiming to better reflect hidden relations not captured by an initially constructed structure. Kim et al. [han] introduced the Hypergraph Attention Network (HAN). HAN constructs hypergraphs from symbolic structures and uses co-attention between hyperedges to integrate multimodal representations. Li et al. [hgtn] designed the Hypergraph Transformer Neural Network (HGTN), which explores Transformer-based attention on heterogeneous information networks to capture higher-order semantics and facilitate communication between vertex and hyperedge types.
Hypergraph learning has shown significant advantages in modeling high-order relationships across various domains such as social networks [social, social2], knowledge graphs [kg], recommendation systems [recom, recom2], action recognition [ActRec_tip, ActRec_tip2], and bioinformatics [bio, bio2].
2.2 Hypergraph Neural Networks in Visual Recognition
@@
In recent years, research has gradually begun to explore the incorporation of hypergraphs and hypergraph neural networks into visual recognition tasks [ActRec_tip, ActRec_tip2, 3d_retr_tip]. Feng et al. proposed Hyper-YOLO [hyperyolo], which integrates a hypergraph computation module into the neck component of YOLO [yolo] to achieve cross-level and cross-spatial hypergraph semantic aggregation. Han et al. introduced Vision HGNN [vihgnn], exploring the replacement of Transformer modules with hypergraph convolution modules. Wang et al. proposed HGFormer [hgformer], which establishes many-to-many topological connections at the patch level and utilizes a hypergraph attention mechanism to learn global visual relationships. Chen et al. proposed HI2R [intra_inter], which simultaneously extracts intra-category structural information and inter-category relational information through hypergraphs for fine-grained recognition.
Most recently, Fixelle et al. proposed Hypergraph Vision Transformers (HgVT) [hgvt], a dedicated vision backbone that unifies hypergraph modeling with Transformer-style attention. HgVT dynamically computes a soft vertex-hyperedge adjacency matrix using a sharpened sigmoid. It then derives a hard adjacency mask for sparse vertex self-attention, while retaining the soft adjacency to modulate the vertex-to-hyperedge and hyperedge-to-vertex communication pathways. HgVT further introduces population regularization to shape hyperedge membership density and adopts an expert-based pooling strategy for prediction. Notably, its vertex-hyperedge communication remains differentiable through the soft adjacency matrix.
Although hypergraph-based visual recognition has made progress, many existing approaches still face limitations in representing visual semantics. Traditional methods [vihgnn, hyperyolo] relying on $k$-NN or $\epsilon$-ball suffer from inherent hyperedge redundancy and quadratic computational costs. Recent attention-based methods such as HAN [han] and HGTN [hgtn] have introduced more flexible frameworks, but the process of constructing the hypergraph in these methods remains non-learnable. HgVT [hgvt] further advances the area by combining sparse vertex self-attention with soft vertex-hyperedge communication. In this paper, we explore a different design point: SoftHGNN is a lightweight add-on block that performs vertex $\rightarrow$ hyperedge $\rightarrow$ vertex message passing directly over a continuous participation matrix, without introducing a separate thresholded vertex-to-vertex self-attention branch. In addition, we scale up the hyperedge capacity efficiently through sparse hyperedge selection with load-balancing regularization.
@@ 3.1 Hypergraph
A hypergraph is an extension of a graph that explicitly models high-order relationships within data. Formally, a hypergraph is represented as: $\mathcal{G}=(\mathcal{V},\mathcal{E}),$ where $\mathcal{V}$ denotes the set of vertices, and $\mathcal{E}$ denotes the set of hyperedges. Unlike traditional graphs, each hyperedge in a hypergraph can simultaneously connect an arbitrary number of vertices. As a result, hypergraphs naturally capture complex high-order associations among multiple vertices. The structure of a hypergraph can be described using an incidence matrix ${H}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{E}|}$, where $H_{v,e}=1$ if $v\in e$ and $H_{v,e}=0$ otherwise.
@@ 3.2 Hypergraph Neural Networks
Despite the theoretical promise of HGNNs in capturing high-order semantic associations, several inherent limitations arise when applying them to visual recognition tasks. First, traditional hypergraph constructions typically rely on methods such as $k$-Nearest Neighbors or $\epsilon$-ball criteria, which generate one hyperedge per vertex. In visual applications, where the number of vertices is extremely large, this approach leads to an explosion of hyperedge numbers. This redundancy not only introduces substantial computational overhead but also results in many hyperedges that capture overlapping or less task-relevant relationships, thereby diluting the model’s focus on truly meaningful associations. Second, the incidence matrix $H$ in the traditional hypergraph enforces a hard binary assignment: a vertex is either fully included in a hyperedge or not at all. Such a rigid formulation fails to capture the inherent continuity and ambiguity present in visual data, especially in scenarios with partial occlusions or overlapping semantic regions, thus limiting the expressiveness and robustness of the model.
@@ 3.3.1 Complexity of Hypergraph Construction
Table 1: Important symbols used in the Proposed Method section.
Symbol Description
$G$ soft hypergraph
$V=\{v_{i}\}_{i=1}^{N}$ vertex set
$E=\{e_{m}\}_{m=1}^{M}$ soft hyperedge set (each $e_{m}$ conceptually involves all vertices)
NN number of vertices in a sample
$M$ preset number of soft hyperedges
$D$ vertex feature dimension
$X\in\mathbb{R}^{N\times D}$ vertex feature matrix; row $i$ is $x_{i}$
$x_{i}\in\mathbb{R}^{D}$ feature of vertex $v_{i}$
$A\in[0,1]^{N\times M}$ participation matrix; $A_{i,m}$ is participation weight of $v_{i}$ in $e_{m}$
$H\in\{0,1\}^{N\times M}$ binary incidence matrix in traditional hypergraph
$P_{0}\in\mathbb{R}^{M\times D}$ globally shared learnable hyperedge prototype parameters
$\Delta P\in\mathbb{R}^{M\times D}$ sample-specific prototype offset
$P\in\mathbb{R}^{M\times D}$ dynamic hyperedge prototype (sample-specific)
$S\in\mathbb{R}^{N\times M}$ unnormalized prototype-vertex feature similarity score.
$\mathcal{E}_{\text{fixed}},\mathcal{E}_{\text{dyn}}$ fixed and dynamic soft hyperedge subsets (Eq. (18))
$M_{\text{fixed}},M_{\text{dyn}}$ sizes of $E_{\text{fixed}}$ and $E_{\text{dyn}}$; $M=M_{\text{fixed}}+M_{\text{dyn}}$
$S_{\text{dyn}}\in\mathbb{R}^{N\times M_{\text{dyn}}}$ columns of $S$ corresponding to dynamic soft hyperedges
$g_{j}$ activation score of dynamic hyperedge $j$ (Eq. (19))
$k$ number of selected dynamic hyperedges (top-$k$)
$\mathcal{E}_{\text{sel}}$ selected dynamic hyperedge subset, $|\mathcal{E}_{\text{sel}}|=k$
$S^{\prime}\in\mathbb{R}^{N\times(M_{\text{fixed}}+k)}$ updated unnormalized assignments after selection
$A_{\text{total}}\in[0,1]^{N\times(M_{\text{fixed}}+k)}$ final participation matrix after selection and Softmax (Eq. (20))
$T$ number of forward passes used to compute activation probabilities
$\mathcal{M}\in\{0,1\}^{T\times M_{\text{dyn}}}$ binary selection mask over dynamic hyperedges across passes (Eq. (21))
$p_{j}\in[0,1]$ activation probability of dynamic hyperedge $j$ (Eq. (22))
$p_{\text{target}}\in[0,1]$ target activation probability.
$\mathcal{L}_{\text{LB}}$ load-balancing regularization loss (Eq. (23))